# Quasi-one-dimensional spin transport in altermagnetic $Z^{3}$ nodal net metals

### v3

@@ Abstract
with $\bm{s}$ the Pauli matrixes acting on spin space and $J_{z}$ denoting the strength of the AM potential. When the AM potential ($J_{z}>0$) is much stronger than other hopping and energy terms, the eight bands are divided into two groups with each group including four bands. Since the two groups are separated in energy space by $2J_{z}$, the model (16) can be simplified as two effectively four-band model [66]. The basis of the upper subsystem is $(|A\uparrow\rangle,|B\uparrow\rangle,|C\downarrow\rangle,|D\downarrow\rangle)^{T}$ and that for the lower subsystem is $(|A\downarrow\rangle,|B\downarrow\rangle,|C\uparrow\rangle,|D\uparrow\rangle)^{T}$. We focus on the upper subsystem and omit $J_{z}$ from it, as it only shifts the bands. Following the classical method [67], the upper four-band model can be expressed as
@@
The band structure of ${\cal H}_{\rm{eff}}$ is calculated and plotted in Ref. [62], which qualitatively reproduces the band structure of $\beta$-Fe2(PO4)O. Our simple four-band effective model hence can serve as a good starting point for the further study of the physics of AM $Z^{3}$ nodal net.
Conclusions.–In conclusion, we have demonstrated the feasibility of achieving the ideal AM $Z^{3}$ nodal net in real materials and shown that the novel Q1D spin transport phenomena inherent to this topological structure. We have detailed the mechanisms whereby electrons with up-spin predominantly flow in one direction, while those with down-spin move primarily in a perpendicular direction, a characteristic distinctly contrasting with conventional 3D bulk materials and Q1D chain materials. This behavior could be effectively detected through the measurement of anisotropic spin resistance, offering a practical approach to exploring the unique physics of AM systems. Our work not only provides a solid foundation for exploring the novel physics of AM systems but also holds great promise for applications in topological spintronics.

### v4

@@ Abstract
with $\bm{s}$ the Pauli matrixes acting on spin space and $J_{z}$ denoting the strength of the AM potential. When the AM potential ($J_{z}>0$) is much stronger than other hopping and energy terms, the eight bands are divided into two groups with each group including four bands. Since the two groups are separated in energy space by $2J_{z}$, the model (8) can be simplified as two effectively four-band model [62]. The basis of the upper subsystem is $(|A\uparrow\rangle,|B\uparrow\rangle,|C\downarrow\rangle,|D\downarrow\rangle)^{T}$ and that for the lower subsystem is $(|A\downarrow\rangle,|B\downarrow\rangle,|C\uparrow\rangle,|D\uparrow\rangle)^{T}$. We focus on the upper subsystem and omit $J_{z}$ from it, as it only shifts the bands. Then the upper four-band model can be expressed as [63]
@@
The band structure of ${\cal H}_{\rm{eff}}$ is calculated and plotted in Ref. [57], which qualitatively reproduces the band structure and the AM $Z^{3}$ nodal net of the $\beta$-Fe2(PO4)O, as well as the Q1D spin transport. Moreover, we find that the Q1D spin transport will disappear when the AM $Z^{3}$ nodal net is removed [57]. Our simple four-band effective model not only further demonstrate the closed relationship between AM $Z^{3}$ nodal net and Q1D spin transport, but also can serve as a good starting point for the subsequent study of the physics of AM $Z^{3}$ nodal net.
Conclusions.–We have demonstrated the feasibility of achieving ideal AM $Z^{3}$ nodal net in real materials and shown that the novel Q1D spin transport phenomena inherent to this topological structure. We have detailed the mechanisms whereby electrons with up-spin predominantly flow in one direction, while those with down-spin move primarily in a perpendicular direction, a characteristic distinctly contrasting with conventional 3D bulk materials and Q1D chain materials. This behavior could be effectively detected through the measurement of anisotropic spin resistance, offering a practical approach to exploring the unique physics of AM systems. Our work not only provides a solid foundation for exploring the novel physics of AM systems but also holds great promise for applications in topological spintronics.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grants Nos. 12004035, 12274112, 12347117 and 12304188) and the China Postdoctoral Science Foundation (Grants Nos. 2024M754079 and GZC20242181). X. Zhang also thanks the sponsorship from S&$\&T$ Program of Hebei (225676163GH). Appendix on numerical methods.– In this work, the electronic conductivity $\sigma(\mu,T)$ was calculated using the Boltzmann transport equation [64]. The conductivities for the two spin channels have a strong dependence on the chemical potential $\mu$ and temperature $T$, which can be calculated as follows [65, 66]:
[…]
where $f(E,\mu,T)$ is the usual Fermi-Dirac distribution function, defined as
$f(E,\mu,T)=\frac{1}{e^{(E-\mu)/k_{B}T}+1},$ (18)
and $\Sigma_{ij}(\epsilon)$ is the transport distribution function tensor, expressed as
$\Sigma_{ij}(\epsilon)=\frac{1}{N_{k}\Omega_{c}}\sum_{n,\mathbf{k}}v_{i}(n,\mathbf{k})v_{j}(n,\mathbf{k})\tau(n,\mathbf{k})\delta(E-E_{n,\mathbf{k}}),$ (19)
where $n$ is the band index, $E_{n,\mathbf{k}}$ is energy of the $n$-th band at momentum point $\mathbf{k}$, $v_{i}(n,\mathbf{k})$ denotes the $i$-th component of the velocity of the $n$-th band at $\mathbf{k}$, $\delta$ stands for Dirac’s delta function, and $\tau(n,\mathbf{k})$ is the relaxation time. $N_{k}$ and $\Omega_{c}$ represent the number of $k$-points used to sample the Brillouin zone and the unit cell volume, respectively. In this work, the relaxation time approximation was adopted [66, 67, 68], in such case $\tau$ is independent of $n$ and $\mathbf{k}$.
For $\beta$-Fe2(PO4)O, the electronic conductivity was calculated at a temperature of 50 K, which is much lower than Néel temperature of the system $T_{N}$ = 408 K, using a $100\times 100\times 100$ momentum mesh via the Wannier90 code [69, 70]. Since the relaxation time of the typical AM metal RuO2 is theoretically estimated as $\sim$ 600 fs [41, 71], we set the relaxation time of $\beta$-Fe2(PO4)O as a constant of 500 fs, which is comparable with that of RuO2. Similarly, for LiTi2O4 and Co2(PO4)O compounds, $T=50$ K, $\tau=500$ fs, and a momentum mesh of $100\times 100\times 100$ are used to calculate the electronic conductivity.